# Learning 3D Object Spatial Relationships from Pre-trained 2D Diffusion Models

### v1

@@ Appendix D Generality of Our methods
Fig. 16 shows that our OOR diffusion still produces plausible outputs even for text prompts that are not seen during training. In the first example, the spatial relation “pour” is learned, but the object categories “moka pot” and “mug” are not seen during training. In the second example, both “steak” and “knife” are categories seen during training, but the spatial relationship of “cutting steak with a knife” is not learned. The last example shows a case of multi-object OOR. The spatial relation of placing a “cutting board” somewhere is not seen during training, but thanks to the generality of OOR diffusion, it is correctly placed on the “table”.

### v2

@@ Appendix E Generality of Our methods
Generality for Unseen Text Inputs. Fig. 17 shows that our OOR diffusion still produces plausible outputs even for text prompts that are not seen during training. In the first example, the spatial relation “pour” is learned, but the object categories “moka pot” and “mug” are not seen during training. In the second example, both “steak” and “knife” are categories seen during training, but the spatial relationship of “cutting steak with a knife” is not learned. The last example shows a case of multi-object OOR. The spatial relation of placing a “cutting board” somewhere is not seen during training, but thanks to the generality of OOR diffusion, it is correctly placed on the “table”.
Figure 18: Scene Editing Results on TRUMANS, OAKINK2 and OmniObject3D. Our scene editing algorithm works on a variety of real and synthetic datasets.
Generality for 3D Scene Editing. Since our OOR can be applied regardless of objects’ textures or shape details, the synthetic-to-real gap is minimal. Fig. 18 demonstrates the effectiveness of our scene editing algorithm on additional datasets which are TRUMANS, OAKINK2, and OmniObject3D, yielding convincing results.